OpenStudy (anonymous):

finding the potential function for a conservative vector field. I have xy dx and (x+y)dy that I'm trying to find a vector valued function of but I can't get the two integrals to work together where the partial derivatives equal those two equations. Any help?

OpenStudy (anonymous):

Don't you mean a potential function for a conservative vector field?

OpenStudy (anonymous):

yes, sorry. my mind is a bit frazzled

OpenStudy (anonymous):

Okay, let's just try some stuff.

OpenStudy (anonymous):

First integration the partial with respect to \(x\).\[ \int xy~dx = \frac{x^2}2y+g(y) \]

OpenStudy (anonymous):

yes thats what I got. and xy+y^2/2 for the other part

OpenStudy (anonymous):

Now we differentiate with respect to \(y\):\[ \frac{x^2}{2}+g'(y) \]Set it equal to partial with respect to \(y\): \[ x+y = \frac{x^2}2+g'(y) \implies g'(y) = x-\frac{x^2}{2}+y \]

OpenStudy (anonymous):

This means it cannot be a potential function, since \(g'\) and \(g\), are not constant with respect to \(x\).

OpenStudy (anonymous):

So there's no solution?

OpenStudy (anonymous):

There is no potential function.

OpenStudy (anonymous):

You can still integrate over the field though.

OpenStudy (anonymous):

Hm... I'm trying to evaluate a line integral and this is the only way my instructor pointed us toward. (She doesn't teach :/ ) I'll look around for anyother way to do it. thanks!

OpenStudy (anonymous):

Another way to test is to use the fact that \(f_{xy}=f_{yx}\). In this case were were getting \(f_{xy}=y\) and \(f_{yx}=1\).

OpenStudy (anonymous):

Do you know what the path is?

OpenStudy (anonymous):

y=x^2

OpenStudy (anonymous):

What are the start and end points?

OpenStudy (anonymous):

(-1,1) and (2,4)

OpenStudy (anonymous):

We can do this.

OpenStudy (anonymous):

Is it the parameterization way? I saw that in Dawkin's notes but I couldn't understand it

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

We will parametrize this curve by letting \(x=t\).

OpenStudy (anonymous):

Then our curve \(\mathbf r\) becomes: \[ \mathbf r(t) = (x,y) = (x,x^2) = (t,t^2) \]

OpenStudy (anonymous):

Our \(d\mathbf r\) becomes: \[ d\mathbf r = \mathbf r'(t)~dt = (1,2t)~dt \]

OpenStudy (anonymous):

Our vector field \(\mathbf F\) becomes: \[ \mathbf F = (xy,x+y) = (tt^2,t+t^2) = (t^3,t+t^2) \]

OpenStudy (anonymous):

wait I was good up until the last one

OpenStudy (anonymous):

oh, thats cuz of x=t. But why is it only the y part of the coordinates?

OpenStudy (anonymous):

All together: \[ \int_C\mathbf F\cdot d\mathbf r = \int_{-1}^2\mathbf F\cdot \mathbf r'(t)~dt = \int_{-1}^2 (t^3,t+t^2)\cdot (t,t^2)~dt \]

OpenStudy (anonymous):

Can you finish it?

OpenStudy (anonymous):

wait, the bounds just changed. Am I supposed to use the X-coord or the Y-coord?

OpenStudy (anonymous):

We could have not done the whole \(x=t\) thing and still arrived at: \[ \int_C\mathbf F\cdot d\mathbf r = \int_{-1}^2\mathbf F\cdot \mathbf r'(x)~dx = \int_{-1}^2 (x^3,x+x^2)\cdot (x,x^2)~dx \]It would be the same.

OpenStudy (anonymous):

and I thought it was F dot ||r'(t)||?

OpenStudy (anonymous):

if that's it then yes, I can finish it

OpenStudy (anonymous):

This is a vector field, not a scalar function.

OpenStudy (anonymous):

You can't dot with a scalar anyway.

OpenStudy (anonymous):

yeah you're right. Alright I think Ive got it

OpenStudy (anonymous):

wait should the last part be (1,2t) instead of (t,t^2)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

okay, so does 15.9167 or 191/12 sound somewhat okay?

OpenStudy (anonymous):

Well, anyway Thanks so much for your help!!!